An electric field of intensity 3.25 kN/C is applied along the x-axis. Calculate the electric flux through a rectangular plane 0.350 m wide and 0.700 m long if the following conditions are true (a) The plane is parallel to the ya-plane (b) The plane is parallel to the xy-plane, (c) The plane contains the y-axis, and its normal makes an angle of 37.0° with the x-axds
Answer #1

formula for electric flux is

flux = EA cos theta

(a)

flux = 3.25* 10^3 N/C * ( 0.350)(0.700) cos 0 = 796.25 N m^2/C

(b)

flux = 3.25* 10^3 N/C * ( 0.350)(0.700) cos 90 = 0

(c)

flux = 3.25* 10^3 N/C * ( 0.350)(0.700) cos 37 = 635.91 N m^2/C